OpenStudy (anonymous):

Simplify 3 fourth root of 112 48 fourth root of 7 2 fourth root of 7 5 fourth root of 7 6 fourth root of 7

OpenStudy (anonymous):

simplify \[3\sqrt[4]{112}\]

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

Mh, did you try anything?

OpenStudy (anonymous):

i don't understand at all.

OpenStudy (the_fizicx99):

Kk, well, you can start by solving 112^1/4, \(\ \sf 3 * \sqrt[4]{112} = \sqrt[4]{2^4 * 7} = \sqrt[4]{2^4} * \sqrt[4]{7} = 2 * \sqrt[4]{7} \) \(\ 3 * 2 * \sqrt[4]{7} = 6\sqrt[4]{7} \)

OpenStudy (anonymous):

so hard :( thank you so much for helping!

OpenStudy (the_fizicx99):

Yw